# Representation and Interpretation in Artificial and Natural Computing

### v1

@@ 6 The Computational Demon and Church Thesis
Church Thesis states that the TM computes the set of computable functions. In the intended sense, a function is computable if it has an effective algorithm: an algorithm that produces the value of an argument if it is defined, or a mark indicating that the function has no value for such an argument, for all the objects in the function’s domain. To know this in general requires knowing whether the machine will halt or not for each of the function’s arguments, and the so-called Halting Machine (HM) would have to be available. It is well known that the HM cannot be a TM (e.g. (2)). Hence, for knowing that a function is computable, it is necessary to find an algorithm that computes it, and show that it is effective in terms of its particular structure. Church Thesis would be refuted if a single function that does not have an effective algorithm were found but computed, as the machine performing such a feat would not be a TM.
@@
Let us define the set $R$ including the demons that compute computable functions. For this, we denominate the functions with finite domain and co-domain, both total and partial, as finite, and all other functions as infinite functions. The finite functions are enumerable through a computable function, and have a unique identifier, which also provides its extensional definition, and can be computed in a table whose columns correspond to its arguments and whose rows to their possible values (13) (Section 11). Hence, all the little demos that compute finite functions are in $R$. The identifiers are the names of the corresponding demons, which can use any particular mode of computing to carry on with the computation, including the table, or an algorithm. That is up to the little demon. The tables help to picture that there are functions that have a clear pattern relating arguments and values systematically, and hence are structured objects, but there are others lacking such a pattern, so the relation between arguments and values is quite arbitrary.
@@
Conversely, there are infinite functions with very little or no structure at all, which can be imaged but are not definable and are non-computable in an absolute sense. The computational demon can achieve the feat of choosing values equally likely among the objects included in an infinite enumerable co-domain, and assign one such value to each object of an infinite enumerable domain; hence, can generate a non-countable number of non-computable functions, and handle them to their corresponding little demons, who are not in $R$. Infinite functions generated by such a procedure cannot be named, defined, or described by people, so we cannot ask the computational demon to compute them, and their corresponding little demos are locked up in a room, a computational limbo, that will remain closed forever.
@@ 7 Natural Computing
If there is natural computing, there should be a natural mode of computing directly above the brain structures involved in cognition and immediately below the knowledge level. The definition of such mode would involve finding the format/s of mental representation/s and the physical mode/s of computing that transforms them. In contrast to artificial computing, in which the entity modifying the representation objectively and the one performing the interpretation subjectively are different, in natural computing both processes are performed by the same agent, so computing and interpreting are two aspects of the same phenomenon, and experiencing the world and being conscious are the manifestations of natural computing. The brain may use different modes for directly transforming representations, and the phenomenological aspect of intentionality may be due to their particular properties. There could even be functions lacking an effective algorithm but having a device used in mental processes, and the brain would be a more powerful computing organ than the Turing Machine. The question of whether there is natural computing is open to empirical investigation, and finding it out would come with solving the hard problem of consciousness. If it is ever found, the representational hypothesis of mind would be upheld. However, if it turns out that such an object does not exist, the non-representational hypothesis holds, and the mind is not a computational process, but something else, not known as yet.
8 Acknowledgments

### v2

@@ 6 The Computational Demon and Church Thesis
Church Thesis states that the TM computes the set of computable functions. In the intended sense, a function is computable if it has an effective algorithm: an algorithm that produces the value of an argument if it is defined, or a mark indicating that the function has no value for such an argument, for all the objects in the function’s domain. To know this in general requires knowing whether the machine will halt or not for each of the function’s arguments, and the so-called Halting Machine (HM) would have to be available. It is well known that the HM cannot be a TM [boolos-jeffrey-1989]. Hence, for knowing that a function is computable, it is necessary to find an algorithm that computes it, and show that it is effective in terms of its particular structure. Church Thesis would be refuted if a single function that does not have an effective algorithm were found but computed, as the machine performing such a feat would not be a TM.
@@
Let us define the set $R$ including the demons that compute computable functions. For this, we denominate the functions with finite domain and co-domain, both total and partial, as finite, and all other functions as infinite functions. The finite functions are enumerable through a computable function, and have a unique identifier, which also provides its extensional definition, and can be computed in a table whose columns correspond to its arguments and whose rows to their possible values [Pineda-2024], Section 11. Hence, all the little demos that compute finite functions are in $R$. The identifiers are the names of the corresponding demons, which can use any particular mode of computing to carry on with the computation, including the table, or an algorithm. That is up to the little demon. The tables help to picture that there are functions that have a clear pattern relating arguments and values systematically, and hence are structured objects, but there are others lacking such a pattern, so the relation between arguments and values is quite arbitrary.
@@
Conversely, there are infinite functions with very little or no structure at all, which can be imaged but are not definable and are non-computable in an absolute sense. The computational demon can achieve the feat of choosing values equally likely among the objects included in an infinite enumerable co-domain, and assign one such value to each object of an infinite enumerable domain; hence, can generate an immeasurable number of non-computable functions, and handle them to their corresponding little demons, who are not in $R$. Infinite functions generated by such a procedure cannot be named, defined, or described by people, so we cannot ask the computational demon to compute them, and their corresponding little demos are locked up in a room, a computational limbo, that will remain closed forever.
@@ 7 Natural Computing
If there is natural computing, there should be a natural mode of computing directly above the brain structures involved in cognition and immediately below the knowledge level. The definition of such mode would involve finding the format/s of mental representation/s and the physical mode/s of computing that transforms them. In contrast to artificial computing, in which the entity modifying the representation objectively and the one performing the interpretation subjectively are different, in natural computing both processes are performed by the same agent, so computing and interpreting are two aspects of the same phenomenon, and experiencing the world and being conscious are the manifestations of natural computing. The brain may use different modes for directly transforming and interpreting representations, and the phenomenological aspect of intentionality may be due to their particular properties. There could even be functions lacking an effective algorithm but having a device used in mental processes, and the brain would be a more powerful computing organ than the Turing Machine. The question of whether there is natural computing is open to empirical investigation, and finding it out would come with solving the hard problem of consciousness. If it is ever found, the representational hypothesis of mind would be upheld. However, if it turns out that such an object does not exist, the non-representational hypothesis holds, and the mind is not a computational process, but something else, not known as yet.
8 Funding statement
@@
9 Conflicts of Interest statement
The author states that there are no conflicts of interest related to this paper.
[…]